OpenStudy (anonymous):

a motorcycle is capable of accelerating at 4.7m/s2. starting from rest, how far can it travel in 3 seconds

OpenStudy (anonymous):

\[d = \frac{ 1 }{ 2 } \times a t ^{2}\] \[d = \frac{ 1 }{ 2 } \times 4.7 \times 3 ^{2}\]\[d = 21.15 m\]